OpenStudy (anonymous):

Please help me with this question. log(x - 1) = -log 2

OpenStudy (mathstudent55):

Remember that a multiplier of the log is an exponent. \(\large a \log b = \log b^a\) Ok?

OpenStudy (anonymous):

yes, i understand.

OpenStudy (mathstudent55):

Then let's start. \(\log(x - 1) = -\log 2\) \(\log(x - 1) = -1 \times\log 2\) \(\log(x - 1) = \log 2^{-1}\) \(x - 1 = 2^{-1} \) \(x - 1 = \dfrac{1}{2} \) \(x = 1 + \dfrac{1}{2} \) \(x = \dfrac{3}{2} \)